OpenStudy (anonymous):

Please help me! Will give medal! :) Darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, sister, and him. Calculate the probability of each compound event.

OpenStudy (anonymous):

Match the term and definition. Probability(Darnell and sister), with replacement Probability(mother and father), without replacement Probability(parent and sister), without replacement Probability(parent and child), with replacement A) one-fourth B) one-sixth C) one-twelfth D) one-sixteenth

OpenStudy (anonymous):

@312856MLP help me please

OpenStudy (anonymous):

P(D,§)=1/4∗1/3=1/12 P(M,F)=1/4∗1/4=1/16 P(P,§)=2/4∗1/4=1/8 P(P,C)=2/4∗2/4=1/4

OpenStudy (anonymous):

there is no 1/8

OpenStudy (anonymous):

@312856MLP thank you.. but I guess you meant 1/6 not 1/8

OpenStudy (anonymous):

yeah and sorry

OpenStudy (anonymous):

Thank you!! :)

OpenStudy (anonymous):

welcome